insert three arthimetic means between -18 and 4
Answer by ikleyn(53763)

.

In the number line, you have this segment  [-18,4]  of the length  18 + 4 = 22 units.


They want you insert 3 numbers (3 points in the number line) to divide this segment in 4 (four) equal parts.


I am sure that you do understand why I am talking about 4 parts: because 4 = (3-1) + 2.


Next, it is clear for you, that each part must be  22/4 = 5.5 units long.


It means that your points (numbers) are the terms of an arithmetic progression
with the first term of -18 and the common difference of 5.5


    -18, -12.5, -7, -1.5, 4.     (*)


So, the three arithmetic means that you insert are the numbers  -12.5, -7, -1.5.


In other words, the terms of this progression  (*), excluding its endpoints, give you the answer to the problem's question.

Solved.